the area of a rectangle is 44 square meters. if the length is one meter less than three times the width, what are the dimensions pf the rectangle? What is the algebraic equation Answer by mananth(16949) 
 Area = 44 sq.ft
Width = x ft
Length = (3x-1 )
Area = x*1(3x-1)
44 =3x^2-x
3x^2-x -44 = 0
3x^2-12x+11x-44=0
3x(x-4)+11(x-4)=0
(x-4)(x+11)=0
x= 4 which is positive
width = 4 ft
Length = 3x-1=> 11 ft
m.ananth@hotmail.ca